Tom deposited $5,544 into a savings account paying 8.2% interest compounded monthly. Find the balance after 8 years.


1
Expert's answer
2021-07-14T10:12:59-0400

"A=P\times(1+\frac{r}{12})^{n\times12}"

"A=5544(1+\frac{0.082}{12})^{8\times12}"

"A=5544(1.9228)"

"A=10659.86"

balance after 8 years is "10659.86dollars"